OpenStudy (anonymous):

factor x² - 16x + 15

OpenStudy (zehanz):

x²-16x+15=(x+a)(x+b). Look for a and b, where a+b=-16 and ab=15

OpenStudy (zehanz):

Hint: they are both negative...

OpenStudy (anonymous):

-1 and -15 are the only one that equal 16

OpenStudy (zehanz):

Yes! -1 + -15 = -16 >>> OK -1 * -15 = 15 >>>OK So x²-16x+15=(x-1)(x-15).

OpenStudy (anonymous):

so I was right

OpenStudy (zehanz):

You were...

OpenStudy (zehanz):

Although you wrote 16 instead of -16 :)